OpenStudy (anonymous):

How can i do this multiplly prime numbers to get 168

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

well you can multpaly but ill take a lot longer unless you did the factor tree

OpenStudy (anonymous):

iam thing of a way to slove it

OpenStudy (anonymous):

Break 168 down into prime factors (ie divide repeatedly by 2, 3, 5, 7,11,....) 168 = 2 x 84 = 2 x 2 x 42 = 2 x2 x2 x2 x21 = 2 x2 x2 x2 x2 x3 x7

OpenStudy (anonymous):

yea thats right now make the \[x ^{?}\]

OpenStudy (anonymous):

the answer is \[2^{5}\] 3,\[7^{1}\]

OpenStudy (anonymous):

i still dn get how to show my work

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

168=2x2x42 = 2x2x2x2x21 = \[2^{5}\] \[3^{1}\] \[7^{1}\]

OpenStudy (anonymous):

that's how you slove the problem

OpenStudy (anonymous):

I'd write \[168 = 2^5 \times 3 \times 7\]

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

bye dude see you tomarow

OpenStudy (anonymous):

bye thx

OpenStudy (anonymous):

gyc could u plz explain yur answer